# Exploring non-Euclidean photonics: Pseudosphere microlaser

### v1

@@ Abstract
Classical and wave properties of pseudosphere-shape microlasers are investigated through experiments and numerical simulations. These pseudosphere microlasers are surface-like organic microlasers with constant negative curvature, which were fabricated with high optical quality by direct laser writing. It is shown that they behave, in many ways, similar to two-dimensional flat disks, regardless of the different Gaussian curvature of the two systems. Indeed, it is evidenced that, due to rotational symmetry, the pseudosphere is an integrable system with marginally stable dynamics.
@@ I Introduction
This first journey in non-Euclidean photonics explores the pseudosphere, a surface with constant negative curvature. Actually surfaces with positive curvature, like the sphere, the bottle [1], or the exterior of a toroid [2] have already given rise to an abundant literature in photonics because of their relative ease of fabrication from glass by melting. On the contrary, negatively curved surfaces - a field of research also called hyperbolic geometry - were intensively investigated by mathematicians [3], theoretical physicists [4, 5, 7, 6], and artists [8], but remained out of reach of experiments until the recent rise of three-dimensionnal (3D) printing at the microscale by Direct Laser Writing (DLW) [9]. This mature technology contributes by now to photonics, as illustrated by the pseudospherical microlaser shown in Fig. 1.
In photonics, negatively curved surfaces provide two main properties:
@@
  • Light no longer propagates along straight lines, but along geodesics, which are the shortest paths between two points.
@@
  • Each geodesic is unstable [10]. That means that any pair of close trajectories will diverge exponentially from each other, which leads to a pure chaos.
@@
Figure 1: (a) SEM image of a tractoid microlaser with radius $R_{0}=20$ µm and thickness $\delta=2$ µm. (b) Schematic of a half tractoid microlaser with the same radius and thickness. The thickness $\delta$ is obtained by vertically translating two tractoid surfaces with respect to each other. The height of the microlaser is around $h=40$ µm
In this paper, which is the first part of our exploration of non-Euclidean photonics, we resort to two main simplifications in order to sharpen our tools before deeper investigations:
@@
  • The curvature is certainly negative, but moreover constant, which allows to analytically calculate the stability of geodesics, as discussed in Sec. IV, and gives access to the Poincaré half-plane (see Sec. II.2), which is a very convenient mapping for studying hyperbolic geometry.
@@
  • We consider therein a surface of revolution with a boundary which is also rotationally invariant. The angular momentum is then conserved, in addition to the energy. As the propagation within the surface is a two-dimensional dynamical system, there are as many constants of motion as degrees of freedom, which leads to integrable dynamics.
@@
Among several surfaces of revolution with constant negative curvature [11], we focus on the tractoid defined in Sec. II. Integrability is tested in Sec. IV where we exhibit the stability ($\equiv$ monodromy) matrices for propagation and reflection and show that each periodic geodesic is indeed marginally stable. The demonstration is similar for classical massive particles and for light rays, as they basically follow the same equations [12].
@@
On the wave-dynamical level (both matter waves or electromagnetic waves), the rotational invariance makes the wave equation separable on the tractoid surface. Then the radial part of the Helmholtz equation (for light) looks like a Schrödinger equation with an effective potential, which can be solved either analytically (Sec. V.2), or numerically (Sec. V.4), showing intriguing similarities to the flat circular disk. Moreover experiments reported in Sec. III evidence lasing emission from Whispering Gallery Modes (WGMs), like in most circular microresonators. Finally Sec. VI concludes the comparison between the flat disk and the tractoid with a full 3D numerical analysis taking into account the actual thickness and polarization.
II Tractoid, Poincaré half-plane and conformal mapping
@@
This section introduces the tractoid surface and one of its representations which is called the Poincaré half-plane.
@@ II.1 Parametrization of the tractoid
Pseudosphere (or pseudospherical surface) is the name given to two-Dimensional (2D) surfaces embedded in the 3D-Euclidean space with constant negative Gaussian curvature. The simplest example of pseudosphere is the tractoid, that is studied hereafter. A tractoid of radius $R_{0}$ is a surface having a Gaussian curvature $-\frac{1}{R_{0}^{2}}$ at each point. For obvious similarities with the sphere of radius $R_{0}$, whose Gaussian curvature is $+\frac{1}{R_{0}^{2}}$, the tractoid is commonly referred to as a pseudosphere in the literature. A typical parametrization of the tractoid surface $\vec{r}=\vec{F}(u,v)$ is:
@@
where $E$ and $G$ are related to the radius of revolution $R=\sqrt{X^{2}+Y^{2}}$ and $Z$ as
@@ II.2 The Poincaré half-plane
which maps the tractoid onto the subspace of $\mathbb{H}$ corresponding to $\mathbb{A}=\{(x,y)\in\mathbb{R}^{2}\mid y\geq 1\}$, preserves both the angles (conformal mapping) and the distances (isometry). Therefore, both the geodesic equation and the wave equation are conserved under this mapping [13]. In particular, this means that distances can be computed on the tractoid as well as in the Poincaré half-plane. For two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$, the length of the geodesic connecting them is [4]
@@
Note that the edge $u=0$ of the tractoid is mapped to the line $y=1$ (blue line in Fig. 2). So the part of the Poincaré plane with $y<1$ cannot be mapped onto the tractoid.
There are several advantages to working in the Poincaré half-plane, like the Cartesian coordinate system, but the main one is that the geodesics have an extremely simple expression. Indeed, there are only two kinds of geodesics in the Poincaré half-plane: segments of half-circles whose centers are on the $x$-axis and, straight vertical lines perpendicular to the $x$-axis, as shown in Fig. 2. Therefore most of our calculations will be done in the Poincaré half-plane.
There exist many other conformal mappings of the tractoid which are not used in this paper. Two are nonetheless discussed in App. A. The first one is the Poincaré disk and the second one is a conformal mapping proposed in [13] which can be applied to any surface of revolution.
@@ III Tractoid microlaser and experimental results
After discussing the fabrication of the tractoid microlasers by Direct Laser Writing (DLW), experimental laser spectra will be analyzed. It will be shown that the lasing modes are essentially WGMs.
@@ III.1 Tractoid fabrication and experimental setup
The 3D design of the tractoid microlasers was created using the free OpenSCAD software by vertically translating two tractoid surfaces of radius $R_{0}=20$ µm and cut at $u_{max}=3$. The translation distance between the two tractoids $\delta=2\,\mu$m defines the thickness of the tractoid microlaser [cf. Fig. 1(b)]. At this point, we should point out the obvious but crucial fact that the tractoid microlaser differs from a mathematical tractoid surface because of its thickness and its finite size.
@@
The tractoid microlasers were pumped individually and with uniform intensity by a beam perpendicular to the substrate from a frequency-doubled Nd:YAG laser (532 nm, 500 ps, 10 Hz). Their emission was analyzed by a spectrometer coupled to a CCD camera, providing an overall resolution of 0.03 nm. The experiments were carried out at room pressure and temperature. The microlaser light was mostly emitted parallel to the substrate plane. The photograph in Fig. 3(b) shows a lasing tractoid microcavity, whose emission mainly originates from the circular boundary at a near grazing angle. This is observed from all directions, which is consistent with the rotational symmetry of the microlaser. Thus, we shall make the hypothesis that the lasing modes of the tractoid microlasers are mainly WGMs, which will be confirmed in the next paragraph. The laser emission spectrum is comb-like, as shown in Fig 4(a). Finally, the tractoid microlaser emission is polarized with the electric field lying parallel to the substrate, which corresponds to Transverse Electric polarization (TE) and will be discussed further in Sec. VI.
Figure 3: Photographs of a tractoid microlaser illuminated with a low intensity white light without pumping (a), and under pumping (b). The green pump light is removed by a notch filter, while the laser emission is orange.
@@ III.2 Spectrum analysis
Figure 4: (a) Experimental laser spectrum of a tractoid microlaser. (b) Normalized Fourier transform of the spectrum in (a) (the original spectrum must be expressed versus wavenumbers instead of wavelengths) using three different numerical methods: (top) Discrete Fourier Transform (DFT) algorithm which assumes that the sampling is regular, while the actual experimental data are not regularly sampled, (middle) same DFT algorithm, but taking into account the actual spacing between data points, and (bottom) Lomb transform. The perimeter of the tractoid microlaser is indicated by the red dashed line.
A typical experimental laser spectrum is presented in Fig. 4(a). It is composed of several frequency combs. If there is only a single frequency comb, the free spectral range or the distance between two adjacent wavenumbers $\Delta k$ is inversely proportional to the geometric length $L$ of the classical periodic trajectory underneath
@@
where $n_{g}$ is the group refractive index. The protocol to determine $n_{g}$ is based on the effective index approximation, which states that the 3D Helmholtz equation can be approximated by the 2D Helmoltz equation within the tractoid surface using the effective index $n_{eff}$, which takes into account the thickness of the tractoid and the bulk index of the resist. This methodology is detailed in App. V of Ref. [9] and yields $n_{g}=1.58\pm 0.05$.
In Fig. 4(b), the Fourier transform of the spectrum was calculated via three different methods, which are discussed in detail in App. B. Our conclusion is that the position of the peaks is more accurately determined with the Lomb method and moreover the result is less noisy.
The Fourier transform is peaked at $n_{g}L=193\pm 6\,\mu$m. This length is slightly smaller than the perimeter of the tractoid $P=n_{g}2\pi R_{0}=198\,\pm 10\,$ µm as shown by the dashed red line in Fig. 4(b). It thus corresponds to WGMs propagating along the circular boundary of the microlaser which is consistent with the emission spots located on the boundary of the tractoid [cf. Fig. 3(b)], as discussed in Sec. III.1.
@@ IV Stability of the periodic geodesics
This section deals with the classification and stability of periodic geodesics on the tractoid. Since the experiments were carried out in the semiclassical regime $kR_{0}\simeq 200$, geometrical optics approach is relevant. Moreover, in the semiclassical regime, the density of states can be expressed as a sum over periodic geodesics with a weighting depending on their stability [14]. We demonstrate here that each periodic geodesic of the full tractoid $u\in[0,+\infty[$ is marginally stable, as expected from the rotational invariance. The case of the truncated tractoid is similar and analyzed in App. C.4
@@ IV.1 Classification of the periodic geodesics
In this subsection, we classify all the periodic geodesics of the full tractoid $u\in[0,+\infty[$. We work in the Poincaré half-plane, since there is a one-to-one correspondence between geodesics on the actual tractoid and in the Poincaré half-plane where their expression is simpler (cf. Sec. II.2).
There exist only two families of geodesics, the vertical lines $x=\,$constant and the arc circles with the center on the $x$-axis $y=0$. A trajectory on the tractoid cannot switch from one kind of geodesics to another. Indeed, the reflection on the boundary at $y=1$ must be normal for the vertical lines, while a geodesic arc circle cannot bounce on the boundary perpendicularly because its center is on the $x$-axis at $y=0$. Therefore, arc circles are the only way to obtain periodic geodesics. At each reflection at $y=1$, the reflection angle is equal to the incident angle. Consequently, a periodic geodesic is a series of identical arc circles with the same radius, as depicted in Fig. 5. Thus, it is uniquely identified by a pair of positive relatively prime integers $(M,P)$ where $M$ is the number of bounces on the boundary and $P$ is the winding number, that is the number of complete rotations of $2\pi$ around the tractoid. In other words, the periodic geodesics on the tractoid have the same structure than periodic orbits of the flat disk.
We can compute the length $L_{M,P}$ of the $(M,P)$-periodic geodesic as well as $\theta_{M,P}$ the angle between the boundary and the geodesic at the reflection (cf. Fig. 5 for the notations):
@@ IV.2 Stability of the periodic geodesics
The stability of the $(M,P)$-periodic geodesic depends on the eigenvalues of its monodromy matrix $\mathcal{M}$, which describes the result of a small perturbation of the initial conditions on the trajectory (see App. C for the definition). Since the $(M,P)$-periodic geodesic corresponds to $M$ repetitions of the same arc circle, $\mathcal{M}$ is the product $\mathcal{M}=(\mathcal{M}_{0})^{M}$ where $\mathcal{M}_{0}=\Pi\,\mathcal{R}$ is the product of $\Pi$ the monodromy matrix for the propagation along a single arc circle and $\mathcal{R}$ the monodromy matrix for reflection on the circular boundary of the tractoid.
After computations detailed in App. C, the monodromy matrix $\Pi(t)$ for the propagation along a geodesic over a distance $t$ is given by
$\Pi(t)=\begin{pmatrix}\cosh\left(\frac{t}{R_{0}}\right)&R_{0}\sinh\left(\frac{t}{R_{0}}\right)\\ \frac{1}{R_{0}}\sinh\left(\frac{t}{R_{0}}\right)&\cosh\left(\frac{t}{R_{0}}\right)\end{pmatrix}$ (15)
@@
where $\epsilon=\pm 1$, and $\kappa_{t}$ is the tangent curvature (also called geodesic curvature) and measures the curvature of the boundary in the plane which is tangent to the surface and is also the plane of incidence of the trajectory. General formulas for $\kappa_{t}$ and $\epsilon$ are given in App. C.2. For a reflection on the lower circular boundary of the tractoid $u=0$, $\epsilon=+1$ and $\kappa_{t}=\frac{1}{R_{0}}$ Then using formulas (13) and (14), one obtains (see App. C.3):
$\operatorname{tr}\left[\mathcal{M}_{0}(M,P)\right]=\operatorname{tr}\left[\Pi(M,P)\mathcal{R}(M,P)\right]=2$ (17)
As $\det\mathcal{M}_{0}=1$, the two eigenvalues of $\mathcal{M}_{0}$ are both equal to 1. Therefore, like for the flat disk, every periodic geodesic is marginally stable. This might come as a surprise because the propagation alone is unstable (ie. $\operatorname{tr}(\Pi)>2$). So it means that the focusing effect from the reflection at the boundary cancels the divergence during propagation. Of course, it is also a consequence of the rotational symmetry of the tractoid.
Finally, using a ray tracing algorithm, we numerically compute the Poincaré Surface of Section (PSOS) of the tractoid. As shown by Fig. 6, every trajectory covers a horizontal line with $\sin\chi=$ constant, which is a manifestation of angular momentum conservation. The phase space structure is identical to that of a flat disk. There are neither stable island nor chaotic sea which is typical for integrable billiard systems.
Figure 6: Poincaré Surface of Section for the truncated tractoid ($u_{max}=2$). Inset: Definition of the incident angle $\chi$ and the curvilinear coordinate $v$ for the position on the boundary.
To conclude this section, every periodic geodesic is marginally stable, as expected for a surface of revolution. It should be noted that the truncated tractoid ($u\in[0,u_{max}]$) is also a surface of revolution. While other periodic geodesics exist in this geometry, their stability is also neutral, as discussed in App. C.4.

### v3

@@ Abstract
Classical and wave properties of microlasers with the shape of a truncated pseudosphere are investigated through experiments and numerical simulations. These pseudosphere microlasers are surface-like organic microlasers with constant negative curvature, which were fabricated with high optical quality by direct laser writing. It is shown that they behave, in many ways, similar to two-dimensional flat disks, regardless of their different Gaussian curvature. We derive the monodromy matrices for geodesics on the pseudosphere and demonstrate that the periodic geodesics are marginally stable. Actually, due to the rotational symmetry, the pseudosphere is an integrable system.
@@ I Introduction
This work on non-Euclidean photonics explores the ray and wave dynamics on the pseudosphere, a surface with constant negative curvature. Actually surfaces with positive curvature, like the sphere, the bottle [1], or the exterior of a toroid [2] have already given rise to an abundant literature in photonics because of their relative ease of fabrication from glass by melting. On the contrary, negatively curved surfaces - a field of research also called hyperbolic geometry - were intensively investigated by mathematicians [3], theoretical physicists [4, 5, 7, 6, 8], and artists [9], while very few experiments were carried out [10, 11] until the recent rise of three-dimensionnal (3D) printing at the microscale by Direct Laser Writing (DLW) [12, 13]. This mature technology contributes by now to photonics, as illustrated by the pseudospherical microlaser shown in Fig. 1.
In photonics, negatively curved surfaces provide two particular properties:
@@
  • On a curved surface light no longer propagates along straight lines, but along geodesics, which are the shortest paths between two points.
@@
  • On a negatively curved surface each geodesic is unstable [14, 15]. That means that any pair of nearby trajectories will diverge exponentially from each other, which is the signature of chaotic dynamics.
@@
Figure 1: (a) SEM image of a tractoid microlaser with radius $R_{0}=20$ µm and thickness $\delta=2$ µm. (b) Schematic of a half tractoid microlaser with the same radius and thickness. The cavity with finite thickness $\delta$ is obtained by vertically shifting two tractoid surfaces with respect to each other. The height of the microlaser is around $h=40$ µm.
In this paper we resort to two simplifications by choosing curved surfaces with the following properties:
@@
  • The curvature is certainly negative, but moreover constant, which allows to analytically calculate the stability of geodesics, as outlined in Sec. IV, and gives access to the Poincaré half-plane (see Sec. II.2), which is a very convenient parametrization for studying hyperbolic geometry.
@@
  • We consider a truncated surface of revolution, which preserves rotational symmetry, implying that the angular momentum is conserved, in addition to the energy. As propagation confined to the surface is a two-dimensional (2D) dynamical system, there are as many constants of motion as degrees of freedom, implying an integrable dynamics.
@@
Among all the possible surfaces with these properties [16], we focus on the tractoid defined in Sec. II. This surface has already been widely investigated, in particular in Ref. [4]. Our main contributions to this field are the stability ($\equiv$ monodromy) matrices for propagation and reflection exhibited in Sec. IV, and derived in App. C. The demonstration is similar for classical massive particles and for light rays, as they basically follow the same equations [17]. We also show that each periodic geodesic is marginally stable, as expected for a classically integrable dynamics.
@@
On the wave-dynamical level (both matter waves or electromagnetic waves), the rotational invariance makes the wave equation separable on the tractoid surface. Then the radial part of the Helmholtz equation (for light) is mathematical equivalent to a Schrödinger equation with an effective potential, which can be solved either analytically (Sec. V.2), or numerically (Sec. V.4), showing intriguing similarities to the flat circular disk. Moreover experiments reported in Sec. III evidence lasing emission from Whispering Gallery Modes (WGMs), like in most circular microresonators. Finally Sec. VI concludes the comparison between the flat disk and the tractoid with a full 3D numerical analysis taking into account the actual thickness and polarization.
II Tractoid and Poincaré half-plane
@@
This first section introduces the tractoid surface and its representation in the Poincaré half-plane.
@@ II.1 Parametrization of the tractoid
A pseudosphere (or pseudospherical surface) is a two-dimensional (2D) surface embedded in the 3D-Euclidean space with constant negative Gaussian curvature. The simplest example of a pseudosphere is the tractoid, which is studied hereafter. A tractoid of radius $R_{0}$ is a surface having a Gaussian curvature $-\frac{1}{R_{0}^{2}}$ at each point. For obvious similarities with the sphere of radius $R_{0}$, whose Gaussian curvature is $+\frac{1}{R_{0}^{2}}$, the tractoid is commonly referred to as a pseudosphere in the literature. A typical parametrization of the tractoid surface $\vec{r}=\vec{F}(u,v)$ is:
@@
where $E$ and $G$ are related to the radius of revolution $R=\sqrt{X^{2}+Y^{2}}$ and to $Z$ as
@@ II.2 The Poincaré half-plane
which maps the tractoid onto the subspace $\mathbb{A}=\{(x,y)\in\mathbb{R}^{2}\mid y\geq 1\}$ of $\mathbb{H}$, preserves both the angles (conformal mapping) and the distances (isometry). Therefore, both the geodesic equation and the wave equation are conserved under this mapping [18]. In particular, this means that distances can be computed on the tractoid as well as in the Poincaré half-plane. The length of a geodesic starting at $(x_{1},y_{1})$ and ending at $(x_{2},y_{2})$ equals [4]
@@
Note that the edge $u=0$ of the tractoid is mapped to the line $y=1$ (blue line in Fig. 2), and only the part of the Poincaré plane above $y=1$ is accessed by the transformation Eq. (II.2).
The main advantage of working in the Poincaré half-plane is that the functional dependence of the geodesics on the coordinates is simple. Indeed, there are only two kinds of geodesics. These are either parts of semicircles, of which the centers are on the x-axis, or they are straight lines $x=const.$ perpendicular to the x-axis, as shown in Fig. 2. Therefore most of our calculations will be done in the Poincaré half-plane.
@@ III Tractoid microlaser and experimental results
In this section we briefly explain the fabrication of the tractoid microlasers by Direct Laser Writing (DLW), the experimental setup, and exhibit results regarding the analysis of experimental laser spectra.
@@ III.1 Tractoid fabrication and experimental setup
In the experiment, we designed a curved waveguide in the shape of the tractoid by shifting two tractoid surfaces of radius $R_{0}=20$ µm with respect to each other. The distance between the two tractoid surfaces, $\delta=2\,\mu$m, defines the thickness of the tractoid microcavity [cf. Fig. 1(b)]. It was cut at $u_{max}=3$ which corresponds to a height of about 40 $\mu$m. The 3D design of the tractoid microlasers was created using the free OpenSCAD. We would like to stress at this point that the tractoid microcavity differs from a mathematical tractoid surface because of its finite size, which is accounted for in the computations by considering truncated tractoids. Another difference is the finite thickness of this and other curved microcavities. Yet, they can still be described by the scalar Helmholtz equation using the concept of the effective index of refraction [13], as outlined below.
@@
The tractoid microlasers were pumped individually and with uniform intensity by a beam perpendicular to the substrate from a frequency-doubled Nd:YAG laser (532 nm, 500 ps, 10 Hz). Their emission was analyzed by a spectrometer coupled to a CCD camera, providing an overall resolution of 0.03 nm. The experiments were carried out at room pressure and temperature. The microlaser light was mostly emitted parallel to the substrate plane. The photograph in Fig. 3(b) shows a tractoid microcavity under pumping. The laser emission mainly originates from the circular boundary at a near grazing angle and is observed from all directions, which is consistent with the rotational symmetry of the microcavity. These observations are reproducible from sample to sample, thusindicating that the lasing modes are mainly WGMs as confirmed by the analysis of the laser emission spectra presented in the next section. A typical comblike laser spectrum is shown in Fig 4(a).
The threshold curve plotted in Fig. 3(c) evidences the laser effect. The laser threshold is about 0.12 MW.cm-2, which is one order of magnitude lower than for microlasers of similar size and materials, evidencing the high quality factor of WGM.
Moreover the tractoid microlaser emission is polarized with the electric field oriented parallel to the substrate, which corresponds to Transverse Electric polarization (TE) and will be discussed further in Sec. VI.
Figure 3: Photographs of a tractoid microlaser illuminated with a low intensity white light without pumping (a), and with pumping (b). The green pump light is removed by a notch filter, while the laser emission is orange. (c) Intensity of the peak at 601.56 nm versus the pump power. The threshold is about 0.12 MW.cm-2. Above 0.2 MW.cm-2 the curve has an irregular increase (not shown), which is a signature of mode competition.
@@ III.2 Spectrum analysis
Figure 4: (a) Experimental laser spectrum of a tractoid microlaser. (b) Normalized Fourier transform of the spectrum in (a) from wavenumber to optical length using three different numerical methods: (top) Discrete Fourier Transform (DFT) algorithm which assumes that the sampling is regular, while the actual experimental data are not regularly sampled, (middle) same DFT algorithm, but taking into account the actual spacing between data points, and (bottom) Lomb transform. The perimeter of the tractoid microlaser is indicated by the red dashed line.
A typical experimental laser spectrum is presented in Fig. 4(a). It is composed of several frequency combs. For each comb, the free spectral range or distance between two adjacent wavenumbers $\Delta k$ is inversely proportional to the geometric length $L$ of the corresponding classical periodic orbit, as may be inferred from the semiclassical approximation of the spectral density [19]
@@
Here $n_{g}$ is the group refractive index, whose calibration protocol is detailed in App. V of Ref. [13] and yields $n_{g}=1.58\pm 0.05$. The meaning of the bulk, effective, and group indices involved in this study is explained in App. A.
In Fig. 4(b), the Fourier transform of the spectrum was calculated via three different methods, which are discussed in detail in App. B. Our conclusion is that the position of the peak is most accurately determined with the Lomb method and moreover the result is less noisy.
The Fourier transform is peaked at $n_{g}L=(193\pm 6)\,\mu$m. This length is slightly smaller than the perimeter of the tractoid $P=n_{g}2\pi R_{0}=(198\,\pm 10)\,$ µm as shown by the dashed red line in Fig. 4(b). It thus corresponds to WGMs propagating along the circular boundary of the microlaser at $u=0$, which is consistent with the observation of emission spots located on the boundary of the tractoid [cf. Fig. 3(b)], as discussed in Sec. III.1.
@@ IV Stability of the periodic geodesics
This section deals with the classification and stability of periodic geodesics on the tractoid. Since the experiments were carried out in the semiclassical regime $kR_{0}\simeq 200$, geometrical optics approach is relevant. In this regime, the spectral density is well approximated in terms of a sum over periodic geodesics with a weighting depending on their stability [20], and thus experimentally accessible through the Fourier transform of the spectra. We demonstrate here that each periodic geodesic of the full tractoid $u\in[0,+\infty[$ is marginally stable, as expected from the rotational invariance. The case of the truncated tractoid is similar and analyzed in App. C.4
@@ IV.1 Classification of the periodic geodesics
To determine the stability of the periodic geodesics of the full tractoid with $u\in[0,+\infty[$, we work in the Poincaré half-plane, since there is a one-to-one correspondence between geodesics on the actual tractoid and in the Poincaré half-plane where their expression is simpler (cf. Sec. II.2).
There exist only two families of geodesics, the vertical lines $x=\,$constant and arcs of circles with center on the $x$-axis $y=0$. A trajectory on the tractoid cannot switch from one kind of geodesics to the other one. Indeed, when moving along a straight-line geodesic, the ray is reflected back at $y=1$, while a ray evolving on a circular-arc type geodesic circle does not impinge the line $y=1$ perpendicularly. Therefore, arc of circles are the only way to obtain periodic geodesics.
Since the angles before and after reflection are the same, the ray propagates along a periodic array of identical circular arcs. An example is depicted in Fig. 5. Thus, periodic trajectories are uniquely identified by a pair of positive, relatively prime integers $(M,P)$ where $M$ is the number of bounces on the boundary $y=1$ and $P$ is the winding number, that is, the number of complete rotations of $2\pi$ around the tractoid. Thus, the periodic geodesics on the tractoid have the same structure as periodic orbits of the flat disk.
The length $L_{M,P}$ of the $(M,P)$-periodic geodesic as well as $\theta_{M,P}$ the angle between the line $y=1$ and the geodesic at the reflection point read (cf. Fig. 5 for the notations):
@@ IV.2 Stability of the periodic geodesics
The stability of the $(M,P)$-periodic geodesic depends on the eigenvalues of its monodromy matrix $\mathcal{M}$, which describes the result of a small perturbation of the initial conditions on the trajectory (see App. C for the definition). Since the $(M,P)$-periodic geodesic corresponds to $M$ repetitions of the same circular arc, $\mathcal{M}$ is the product $\mathcal{M}=(\mathcal{M}_{0})^{M}$ where $\mathcal{M}_{0}=\mathcal{P}\,\mathcal{R}$ is the product of $\mathcal{P}$ the monodromy matrix for the propagation along a single circular arc, and $\mathcal{R}$ the monodromy matrix for reflection on the circular boundary of the tractoid.
After computations based on Ref. [21] and detailed in App. C, the monodromy matrix $\mathcal{P}(t)$ for the propagation along a geodesic over a distance $t$ is given by
$\mathcal{P}(t)=\begin{pmatrix}\cosh\left(\frac{t}{R_{0}}\right)&R_{0}\sinh\left(\frac{t}{R_{0}}\right)\\ \frac{1}{R_{0}}\sinh\left(\frac{t}{R_{0}}\right)&\cosh\left(\frac{t}{R_{0}}\right)\end{pmatrix}$ (15)
@@
where $\epsilon=\pm 1$, and $\kappa_{t}$ is the modulus of the tangent curvature (also called geodesic curvature), which measures the curvature of the boundary in the plane which is tangent to the surface and is also the plane of incidence of the trajectory. General formulas for $\kappa_{t}$ and $\epsilon$ are given in App. C.2. For a reflection on the lower circular boundary of the tractoid $u=0$, $\epsilon=+1$ and $\kappa_{t}=\frac{1}{R_{0}}$ Then using formulas (13) and (14), one obtains (see App. C.3):
$\operatorname{tr}\left[\mathcal{M}_{0}(M,P)\right]=\operatorname{tr}\left[\mathcal{P}(M,P)\mathcal{R}(M,P)\right]=2$ (17)
As $\det\mathcal{M}_{0}=1$, the two eigenvalues of $\mathcal{M}_{0}$ are both equal to 1. Therefore, like for the flat disk, every periodic geodesic is marginally stable. This might come as a surprise because the propagation itself is unstable, (ie. $\operatorname{tr}(\mathcal{P})>2$. Yet, the focusing effect from the reflection at the boundary cancels the divergence during propagation. Of course, it is also a consequence of the rotational symmetry of the tractoid.
Finally, using a ray tracing algorithm, we numerically compute the Poincaré Surface of Section (PSOS) of the tractoid. As shown by Fig. 6, every trajectory covers a horizontal line with $\sin\chi=$ constant, which is a manifestation of angular momentum conservation. The phase space structure is identical to that of a flat disk. There are neither stable islands nor a chaotic sea which is typical for integrable billiard systems.
Figure 6: Poincaré Surface of Section for the truncated tractoid ($u_{max}=2$). Each color corresponds to a specific initial condition ($v$, $\chi$). Inset: Definition of the incident angle $\chi$ and the curvilinear coordinate $v$ for the position on the boundary.
To conclude this section, every periodic geodesic is marginally stable, as expected for a surface of revolution. It should be noted that the truncated tractoid ($u\in[0,u_{max}]$) is also a surface of revolution. While other periodic geodesics exist in this geometry, their stability is also neutral, as outlined in App. C.4.